OpenStudy (anonymous):

What is 3x^((5)/(7)) in radical form?

OpenStudy (anonymous):

Recall that the numerator of fractional exponents are powers, and the denominators are roots. In this case\[3x^{5/7}=3\sqrt[7]{x^5}=3(\sqrt[7]{x})^5\]